# LLM Compression with Neural Architecture Search

### v1

@@ 1 Introduction
Large language models (LLMs) represent a significant advancement in artificial intelligence and are increasingly deployed in products such as chatbots and coding assistants. However, their substantial parameter count leads to high latency and significant demands for computational resource during inference. This renders deployment in time-sensitive or resource-constrained settings, such as embedded systems, often impractical, or results in elevated costs per user query, for example in web services.
@@
  • Parameter-efficient fine-tuning: Fine-tuning LLMs with more than 7B parameters require access to large-scale compute. Building on previous work (Munoz et al., 2024), we empirically investigate the combination of two-stage NAS with state-of-the-art parameter-efficient fine-tuning methods, such as LoRA (Hu et al., 2021), to scale to larger models.
@@
Section 2 provides an overview of related work on compressing LLMs based on distillation, structural pruning, as well as neural architecture search. Section 3 introduces two stage NAS, and Section 4 provides an overview of our approach to scale weight-sharing based NAS to LLMs. We provide an empirical analysis of our approach across baseline methods from the structural pruning literature and an in-depth analysis of our approach in Section 5.
@@ 3 Background and Notations
This section defines the notations and terminologies used throughout the paper. We begin with an overview of the modular components in transformers in 3.1, followed by an introduction to the various building blocks for neural architecture search in 3.2.
@@ 3.1 Transformer Architecture
The transformer architecture (Vaswani et al., 2017) underpins many leading LLMs. We focus on decoder-only architectures (Radford et al., 2019), prevalent in state-of-the-art models. A transformer comprises of an embedding layer followed by $L$ blocks, each containing a self-attention and a fully connected layer, with layer normalization (Ba et al., 2016) or RMS normalization (Zhang and Sennrich, 2019) applied between layers. We delineate the key components for our search space in Section 4.
@@
Fully Connected Layer: The FFN layer is defined as $FFN({\bm{X}})={\bm{W}}_{fc1}\sigma({\bm{W}}_{fc0}{\bm{X}})$, with ${\bm{W}}_{fc0}\in\mathbb{R}^{U\times d_{model}}$, ${\bm{W}}_{fc1}\in\mathbb{R}^{d_{model}\times U}$, and $U=r\cdot d_{model}$. The function $\sigma(\cdot)$ represents a non-linear activation.
@@ 3.2 Neural Architecture Search
Given a search space $\Theta$ composed of architectural design choices, for example the number of heads or layers, neural architecture search (NAS) (see Elsken et al. (2019) for an overview) finds the optimal architecture $\theta^{\star}\in argmin_{\theta\in\Theta}f(\theta)$ that minimizes some error metric $f(\theta)$, such as the validation error $f(\theta)=\mathcal{L}_{valid}(\theta,w_{\theta}^{\star})$ after training the network $w_{\theta}^{\star}=\operatorname*{arg\,min}\mathcal{L}_{train}(w_{\theta})$ with a set of weights $w_{\theta}\in\mathbb{R}^{n}$. The search space $\Theta$ is usually large but finite. We can extend this to multiple objectives $min_{\theta\in\Theta}\left\{f_{0}(\theta),...,f_{k}(\theta)\right\}$, such as for example validation error, inference latency or parameter count. In the multi-objective setting we do not have a single solution $\theta_{\star}$ that simultaneously optimizes all objectives. Instead, we try to approximate the Pareto Set: $P_{f}=\{\theta\in\Theta|\nexists\theta^{\prime}\in\Theta:\theta^{\prime}\succ\theta\}$ of points that dominate all other points in the search space in at least one objective, where we define $\theta\succ\theta^{\prime}$ iff $f_{i}(\theta)\leq f_{i}(\theta^{\prime}),\forall i\in[k]$ and $\exists i\in[k]:f_{i}(\theta)<f_{i}(\theta^{\prime})$.
Two-stage Neural Architecture Search. Classical NAS approaches propose evolutionary algorithms (Real et al., 2019) or reinforcement learning methods (Zoph and Le, 2017) to tackle this optimization problem. However, since every evaluation of $f(\theta)$ involves training and validating a neural network model, these methods consume a substantial amount of compute making them infeasible in practical settings. Two-stage NAS (Yu et al., 2020) trains a super-network ${\mathbf{w}}_{\Theta}^{\star}=\operatorname*{arg\,min}\mathcal{L}_{train}({\mathbf{w}}_{\Theta})$ with a single set of weights ${\mathbf{w}}_{\Theta}\subset\mathbb{R}^{d}$ The super-network is defined such that it contains all possible networks $\theta\in\Theta$ in the search space. After training the super-network, we can compute the validation error $f(\theta)=\mathcal{L}_{valid}(\theta,\hat{{\mathbf{w}}}_{\theta})$ of an architecture $\theta$ by a subset of the weights of the super-network $\hat{{\mathbf{w}}}_{\theta}\subseteq{\mathbf{w}}_{\Theta}$, reducing the overall compute by orders of magnitude. We refer to a network described by $\theta$ that uses a subset of the super-network weights as sub-network.
@@ 4.1 Sampling Distribution
Figure 2: Parameter count of sub-networks a) sampled uniformly at random from $\Theta$, and b) sampled according our fix grid sampling process (described in the main text) for Llama 3.1-8B model. We can see that sampling uniformly oversamples tiny models that are not capable to achieve reasonable performance. Figure c) depicts the calibration procedure and d) represents sub-network perplexities, before and after importance sorting
@@ 4.2 Importance scoring
If we select sub-networks out of the super-network we always select the first elements. For example, remember that $\Theta=\Theta_{d}\times\Theta_{H}\times\Theta_{r}\times\Theta_{L}$, for $\theta=[768,8,2,9]$ we select the first $d_{model}=768$ entries of the embedding vector, the first $H=8$ heads of the multi-head attention layers, the first $U=d_{model}*2$ units in the FFN layer and the first $L=9$ layers. This convention ensures a bijective mapping from $\Theta$ to sub-networks  (Klein et al., 2024), however it reduces the flexibility of selecting sub-networks.
Transformers are invariant to a fixed permutation of neurons throughout the network. Leveraging this, we rearrange the different components of the pre-trained super-network based on their importance score. Following (Muralidharan et al., 2024), we compute for each component the feature importance score on a small calibration dataset, for example Wikitext-2 test set. Figure 2(d) show how this simple permutation significantly reduces subnetwork perplexity, leading to better initialization for super-network training (Section 3.2).
4.3 Combining NAS with Parameter Efficient Fine-Tuning
@@ 5 Experiments
For the empirical evaluation of our model we consider Pythia-410M, Pythia-1B, and Pythia-2.8B from the Pythia family (Biderman et al., 2023), LLaMA-2-7B from the LLaMA-2 family (Touvron et al., 2023), along with the newly introduced LLaMA-3.1-8B. We evaluate sub-networks on five commonsense reasoning tasks in a zero-shot manner using LM-Eval-Harness (Gao et al., 2024). We fine-tune super-networks (see Section 4.2) on the Alpaca (Taori et al., 2023) dataset.
@@ 5.1 Comparison to Structural Pruning Baselines
We compare against the following structural pruning baselines: SliceGPT (Ashkboos et al., 2024) is a post-training structural pruning technique based on Principal Component Analysis that reduces embedding dimensions by removing rows and columns from weight matrices. N:M (Zhou et al., ) prunes N out of every M consecutive weights based on their magnitude. SparseGPT (Frantar and Alistarh, 2023b) prunes weights iteratively to a target sparsity by using a pre-computed Hessian to quantify the sensitivity of each neuron. Wanda (Sun et al., 2024) similar to SparseGPT but uses a simpler pruning metric based on weight magnitude and input activation norms.

### v2

# Large Language Model Compression with Neural Architecture Search

@@ 1 Introduction
Large language models (LLMs) represent a significant advancement in artificial intelligence and are increasingly deployed in products such as chatbots and coding assistants. However, their substantial parameter count leads to high latency and significant computational demands during inference. This renders deployment in time-sensitive or resource-constrained settings, such as embedded systems, often impractical, or results in elevated costs per user query, for example in web services.
@@
  • Parameter-efficient fine-tuning: Fine-tuning LLMs with more than a billion parameters require access to large-scale compute. Building on previous work (Munoz et al., 2024), we empirically investigate the combination of two-stage NAS with state-of-the-art parameter-efficient fine-tuning methods, such as LoRA (Hu et al., 2021), to scale to larger models.
@@
Section 2 provides an overview of related work on compressing LLMs based on distillation, structural pruning, as well as NAS. Section 3 introduces two stage NAS, and Section 4 provides an overview of our approach to scale weight-sharing based NAS to LLMs. We provide an empirical analysis of our approach across baseline methods from the structural pruning literature and an in-depth analysis of our approach in Section 5.
@@ 3 Background and Notations
This section defines the notations and terminologies used throughout the paper. We begin with an overview of the modular components in transformers in 3.1, followed by an introduction to the various building blocks for NAS in 3.2.
@@ 3.1 Transformer Architecture
The transformer architecture (Vaswani et al., 2017) underpins many leading LLMs. We focus on decoder-only architectures (Radford et al., 2019), prevalent in state-of-the-art models. A transformer comprises of an embedding layer followed by $L$ blocks, each containing a self-attention and a fully connected layer, with layer normalization (Ba et al., 2016) or RMS normalization (Zhang and Sennrich, 2019) between layers. We delineate the key components for our search space in Section 4.
@@
Fully Connected Layer: The FFN layer is defined as $FFN({\bm{X}})={\bm{W}}_{fc1}\sigma({\bm{W}}_{fc0}{\bm{X}})$, with ${\bm{W}}_{fc0}\in\mathbb{R}^{U\times d_{model}}$, ${\bm{W}}_{fc1}\in\mathbb{R}^{d_{model}\times U}$, and $U=r\cdot d_{model}$. $\sigma(\cdot)$ represents a non-linear activation function.
@@ 3.2 Neural Architecture Search
Given a search space $\Theta$ composed of architectural design choices, for example the number of heads or layers, NAS (see Elsken et al. (2019) for an overview) finds the optimal architecture $\theta^{\star}\in argmin_{\theta\in\Theta}f(\theta)$ that minimizes some error metric $f(\theta)$, such as the validation error $f(\theta)=\mathcal{L}_{valid}(\theta,w_{\theta}^{\star})$ after training the network $w_{\theta}^{\star}=\operatorname*{arg\,min}\mathcal{L}_{train}(w_{\theta})$ with a set of weights $w_{\theta}\in\mathbb{R}^{n}$. The search space $\Theta$ is usually large but finite. We can extend this to multiple objectives $min_{\theta\in\Theta}\left\{f_{0}(\theta),...,f_{k}(\theta)\right\}$, such as for example validation error, inference latency or parameter count. In the multi-objective setting we do not have a single solution $\theta_{\star}$ that simultaneously optimizes all objectives. Instead, we try to approximate the Pareto Set: $P_{f}=\{\theta\in\Theta|\nexists\theta^{\prime}\in\Theta:\theta^{\prime}\succ\theta\}$ of points that dominate all other points in the search space in at least one objective, where we define $\theta\succ\theta^{\prime}$ iff $f_{i}(\theta)\leq f_{i}(\theta^{\prime}),\forall i\in[k]$ and $\exists i\in[k]:f_{i}(\theta)<f_{i}(\theta^{\prime})$.
Two-stage NAS. Classical NAS approaches propose evolutionary algorithms (Real et al., 2019) or reinforcement learning methods (Zoph and Le, 2017) to tackle this optimization problem. However, since every evaluation of $f(\theta)$ involves training and validating a neural network model, these methods consume a substantial amount of compute making them infeasible in practical settings. Two-stage NAS (Yu et al., 2020) trains a super-network ${\mathbf{w}}_{\Theta}^{\star}=\operatorname*{arg\,min}\mathcal{L}_{train}({\mathbf{w}}_{\Theta})$ with a single set of weights ${\mathbf{w}}_{\Theta}\subset\mathbb{R}^{d}$ The super-network is defined such that it contains all possible networks $\theta\in\Theta$ in the search space. After training the super-network, we can compute the validation error $f(\theta)=\mathcal{L}_{valid}(\theta,\hat{{\mathbf{w}}}_{\theta})$ of an architecture $\theta$ by a subset of the weights of the super-network $\hat{{\mathbf{w}}}_{\theta}\subseteq{\mathbf{w}}_{\Theta}$, reducing the overall compute by orders of magnitude. We refer to a network described by $\theta$ that uses a subset of the super-network weights as sub-network.
@@ 4.1 Sampling Distribution
Figure 2: Parameter count of sub-networks sampled from Llama 3.1-8B model. a) Uniform random sampling from $\Theta$, which oversamples tiny models and b) Grid sampling, which samples more uniformly from $\Theta$
Figure 3: c) Calibration procedure based on architecture magnitude for subnetworks, and d) sub-network perplexities before and after applying importance sorting to the model.
@@ 4.2 Importance scoring
If we select sub-networks from the super-network we always select the first few neorons or layers. For example, remember that $\Theta=\Theta_{d}\times\Theta_{H}\times\Theta_{r}\times\Theta_{L}$, for $\theta=[768,8,2,9]$ we select the first $d_{model}=768$ entries of the embedding vector, the first $H=8$ heads of the multi-head attention layers, the first $U=d_{model}*2$ units in the FFN layer and the first $L=9$ layers. This convention ensures a bijective mapping from $\Theta$ to sub-networks  (Klein et al., 2024), however it reduces the flexibility of selecting sub-networks.
Transformers are invariant to a fixed permutation of neurons throughout the network. Leveraging this, we rearrange the different components of the pre-trained super-network based on their importance score. Following (Muralidharan et al., 2024), we compute for each component the feature importance score on a small calibration dataset, for example Wikitext-2 test set. Figure 3(b) show how this simple permutation significantly reduces subnetwork perplexity, leading to better initialization for super-network training (Section 3.2).
4.3 Parameter Efficient Fine-Tuning for Two-stage NAS
@@ 5 Experiments
For the empirical evaluation of our model we consider Pythia-410M, Pythia-1B, and Pythia-2.8B from the Pythia family (Biderman et al., 2023), LLaMA-2-7B from the LLaMA-2 family (Touvron et al., 2023), along with the newly introduced LLaMA-3.1-8B as well as Phi-2 and Phi-3 (phi3). We evaluate sub-networks on five commonsense reasoning tasks in a zero-shot manner using LM-Eval-Harness (Gao et al., 2024). We fine-tune super-networks (see Section 4.2) on the Alpaca (Taori et al., 2023) dataset.
@@ 5.1 Comparison to Structural Pruning Baselines
We compare against the following structural pruning baselines on the Phi-2, Phi-3, Llama-2-7B and the Llama-3.1-8B model: SliceGPT (Ashkboos et al., 2024) is a post-training structural pruning technique based on Principal Component Analysis that reduces embedding dimensions by removing rows and columns from weight matrices. N:M (Zhou et al., ) prunes N out of every M consecutive weights based on their magnitude. SparseGPT (Frantar and Alistarh, 2023b) prunes weights iteratively to a target sparsity by using a pre-computed Hessian to quantify the sensitivity of each neuron. Wanda (Sun et al., 2024) similar to SparseGPT but uses a simpler pruning metric based on weight magnitude and input activation norms.
@@ 6 Discussion and Conclusion
Compared to classical structural pruning approaches, NAS provides a Pareto set of sub-networks to allow practitioners to select the optimal sub-network matching their constraints. However, compared to some structural pruning methods, NAS requires an additional fine-tuning step to account for training the super-network.
@@ Acknowledgments
Appendix A Extended Related Work
Model compression reduces the size or computational complexity of a neural network model while minimizing loss in performance. It involves techniques such as quantization (reducing the precision of weights and activations), pruning (removing neurons or connections from a model), knowledge distillation (transferring knowledge from a large to a small network e.g. in the form of representations or outputs), low-rank factorization, efficient model design and NAS (see e.g. zhu-arxiv2023 for an overview). In the following, we focus on pruning and NAS.
Pruning removes weights and connections of a network to reduce the number of parameters and accelerate inference. Unstructured pruning removes arbitrary weights while structural pruning considers entire groups of parameters such as attention heads michel-nips19 or layers [sajjad-2023] for removal which is better suited for model acceleration on hardware optimized for dense computation [mishra-2021]. Pruning and in particular structured pruning approaches can result in a loss of accuracy and most pruning methods include a retraining phase to recover as much accuracy as possible. Recent work focused on pruning LLMs to tackle the particular challenges that come with their large number of parameters, the high computational complexity, and the often limited availability of data for retraining. Methods such as ShortGPT [men-arxiv2024] and LaCo [yang-arxiv2024] use importance scores to prune or merge layers of LLMs. SparseGPT [Frantar and Alistarh, 2023a] approximates the optimal weights in a pruning mask using a row-wise iterative update scheme to do unstructured and semi-structure pruning of generative pretrained transformers. Wanda [Sun et al., 2024] extends magnitude pruning [han-neurips2015] by including the activation values on a small calibration set to do unstructured and N:M structured pruning. Flextron [cai-arxiv2024] propose a procedure that allows to extract models for different deployment scenarios by first making by combining an elastic model (cf. Cai et al. [2020]) with methods from mixture of experts (see e.g. fedus-arxiv2022 for a recent review). The router networks can take static information such as a target latency into account but also input-adaptive routing. Probably the closest work to our approach is Minitron [Muralidharan et al., 2024], which uses activation-based importance scores to prune models and knowledge distillation from an uncompressed teacher for retraining. However, they need to considerably reduce the number of architectures that are compared to reduce training time. In our work, we consider much larger search spaces and leverage one-shot NAS for efficient training including knowledge distillation and incorporate importance scores during the architecture sampling procedure. This allows us to combine everything into a single one-step training procedure and calculate the full Pareto front instead of single architectures.
Neural Architecture Search automates the design of deep neural networks in a data-driven manner (see e.g. Elsken et al. [2019], White et al. [2023] for an overview). NAS has been extended to a multi-optimization problem taking also efficiency on a target hardware platform into account such as latency or energy consumption [elsken-iclr19, Cai et al., 2020, Wang et al., 2020] making it closely related to model compression. To tackle the enormous computational cost of early NAS methods [Zoph and Le, 2017, Real et al., 2019], weight-sharing based NAS [saxena-neurips2016, bender-icml18] trains a single super-network from which the weights can be inherited to all architectures in a search space for performance evaluation without further training. A particularly prominent approach to use NAS for model compression is two-stage NAS, which has a dedicated super-network training and multi-objective search stage [bender-icml18, guo-eccv2020, Cai et al., 2020]. Most two-stage methods use the notation of elastic layers [Cai et al., 2020] that can dynamically adjust its size (e.g. width) during training. The training of the supermodel is typically done as proposed in yu-ICCV2019 using the sandwich rule, which aggregates the gradients of multiple sub-networks from the super-network, as well as in-place distillation, which uses the outputs of the largest network in a super-network as targets for smaller ones. Furthermore, tang-iccv2023 and wang-cvpr2021 proposed different strategies how to sample models from supermodel to better cover the Pareto front. A detailed study of NAS for structural pruning has been conducted in Klein et al. [2024] that shows that NAS is a competitive technique to other pruning approaches and highlights in particular the increased flexibility and automation potential of NAS methods that allow to estimate the full Pareto front [Cai et al., 2020, Sukthanker et al., 2024b] instead of having to set a single threshold for pruning. However, even two-stage NAS methods still have a large computational overhead compared to regular model training. To further reduce the computational complexity of NAS, several works proposes to leverage pretrained weights as well as parameter efficient fine-tuning methods. InstaTune [sridhar-iccv2023] uses a pretrained model to initialize and train a super-network on a fine-tuning task. LoNAS [Munoz et al., 2024] freezes the weights of a pretrained backbone and introduces elastic LoRA adapters [Hu et al., 2021]. Similarly, Shears [munoz-arxiv2024] combines unstructured pruning of a pretrained model with NAS to search for elastic LoRA adapters to mitigate the performance loss of pruning.
Appendix B Experimental details
In this section we provide details on the hyperparameters, search spaces and datasets used for fine-tuning different pretrained models studied in the paper. Our hyperparameter settings largely follow litgpt configs 111litgpt config hub.
B.1 Fine-tuning Dataset
We fine-tune all models with the Alpaca[Taori et al., 2023] dataset, which contains 51,759 instruction pairs. Alpaca dataset is generated instruction-following demonstrations by building upon the self-instruct method, which uses an existing strong language model to automatically generate instruction data.
B.2 Pythia
Hyperparameters.
We use a learning rate of 2e-5 for all pythia models and sampling schemes. Furthermore we use a batch size of 4 for pythia-410m and pythia-1b and a batch size of 1 for full fine-tuning of pythia-2.8b. We use the AdamW optimizer with 300 warmup steps and cosine annealing of the learning rate for all models. We fine-tune all the models for 5 epochs.
Search Space.
We define range of choices for search space of pythia models in Table 2, where model_embedding_dim corresponds to the embedding dimension of the pretrained model, the model_number_of_heads corresponds to the number of heads in the pretrained model and model_num_layers corresponds to number of layers in the pretrained model. All other architecture dimensions are same as the respective pretrained models.
Table 2: Configuration Search Space Parameters for Pythia Models
Parameter Type Choice Range [max, min]
[…]
mlp_expansion_ratio Random Integer $[1,4]$
[…]
B.3 Llama
Hyperparameters.
We use a learning rate of 0.0002 for all llama models and sampling schemes. Furthermore we use a batch size of 8 for Llama-3.1-8B and 16 for Llama-2-7B and use the AdamW optimizer with 10 warmup steps with cosine annealing of the learning rate. We fine-tune all the models for 5 epochs. We set the LoRA rank to 32, the LoRA alpha to 16 and the dropout to 0.05. Furthermore, we add LoRA layers to all attention layers, mlp layers and the prediction head.
Search Space.
We define range of choices for search space of Llama-3.1-8B models in Table 3 and for Llama-2-7B in Table 4, where model_embedding_dim corresponds to the embedding dimension of the pretrained model, and model_num_layers corresponds to number of layers in the pretrained model. All other architecture dimensions (including number of heads) are same as the respective pretrained models.
Table 3: Configuration Search Space Parameters for Llama-3.1-8B
Parameter Type Choice Range [max, min]
[…]
mlp_expansion_ratio Random Choice $\{1.0,2.0,3.0,3.5\}$
[…]
Table 4: Configuration Search Space Parameters for Llama-2-7B
Parameter Type Choice Range [max, min]
[…]
mlp_expansion_ratio Random Choice $\{1.0,2.0,2.5,2.6875\}$
[…]
B.4 Phi
Hyperparameters.
We use a learning rate of 0.0002 for all phi models and sampling schemes. Furthermore we use a batch size of 32 and use the AdamW optimizer with 10 warmup steps with cosine annealing of the learning rate. We fine-tune all the models for 1 epoch. We set the LoRA rank to 8, the LoRA alpha to 16 and the dropout to 0.05. Furthermore, we add LoRA layers to all attention layers, mlp layers and the prediction head.
B.5 Search Space.
We define the search space for Phi-2 in a manner similar to Pythia models, Table 2 and define the search space for Phi-3 in Table 5, where model_embedding_dim corresponds to the embedding dimension of the pretrained model, and model_num_layers corresponds to number of layers in the pretrained model. All other architecture dimensions (including number of heads for Phi-3) are same as the respective pretrained models.
Table 5: Configuration Search Space Parameters for Phi-3
Parameter Type Choice Range [max, min]
[…]
mlp_expansion_ratio Random Choice $\{1.0,2.0,2.5,2.6666666666666665\}$
[…]
B.6 Grid Sampling Scheme
We use uniform size of 22 for the architecture grid for Grid-params, Importance-grid-params and Importance-calibrate-mag schemes as defined in 5.2. Furthermore when doing rejection sampling based on parameter count to obtain architectures in different parameter bins, we allow for at most 1000 architecture sampling trials.